# Algebraic geometry of bubbling Kähler metrics

### v1

@@ 2.5. Re-construction of minimal bubble via valuation
Taking Theorem 2.4 ($\mathrm{x}$) and Proposition 2.23 ($\mathrm{ii}$) into account, for instance, it is reasonable to pass our context to general valuations and discuss characterization of the critical valuations in the sense of Theorem 2.4 ($\mathrm{x}$). Our discussions in this subsection are totally coordinates-free.
@@
    For each $v_{X}\in\mathop{\mathrm{Val}}\nolimits^{\rm qm}(x\in X)$, $a\in\mathbb{R}_{>0}$, with $a\ll 1$, say $a\in(0,a_{0})$ for some $a_{0}\in\mathbb{R}_{>0}$, we have a unique valuation $v\in\mathop{\mathrm{Val}}\nolimits^{\rm qm}(x\in\mathcal{X})$ with $\mathcal{J}_{m}(v)|_{X}=J_{m}(v_{X})$ and $v(t)=\frac{1}{a}$.
@@
If $v_{a}(t)=\frac{1}{a}>v_{a}(z_{i})$, then by (8), the independencity of $v_{a}$ with respect to the lifted coordinates $\tilde{z_{i}}$ easily follows. This confirms ($\mathrm{i}$). On the other hand, if we fix a section $\sigma$, then we can impose the condition $\tilde{z_{i}}(\sigma(t))$ for each $i$. Under such constraints, $f_{i}$ of (8) does not have constant terms i.e., positive multiplicities with respect to $\tilde{z_{i}}$s. In that case, if $v_{a}(t)=\frac{1}{a}>\max\{c_{i}\}-\min\{c_{i}\}$, then the obtained quasi-monomial $v_{a}$ does not depend on the lift $\tilde{z_{i}}$s. Hence the proof of ($\mathrm{ii}$). ∎
@@
S.Boucksom kindly suggested to compare the construction with the composite (cf., e.g., [FK18, §6.4]) of the discrete valuation ${\rm ord}_{\mathcal{X}_{0}}$ on $\mathcal{O}_{\mathcal{X},x}$ and $v_{X}$ on $\mathcal{O}_{\mathcal{X}_{0},x}$. Although it gives a height $2$ (not real-valued) valuation, the author sees this as a kind of infinitesimal analogue of the above construction $\{v_{a}\}_{a}$, which may be its integral curve in some unknown sense.
@@
For many ADE singularities with $n=2$, $a_{0}$ can be taken as $\infty$, while [dBS23, Example 7] shows that it is not necessarily the case for $n(>2)$-dimensional $A$-type singularities.
In this paper, we introduce yet another couple of “Yau-Tian-Donaldson type” conjectures for mild degenerations, after [Sun23, dBS23]. It hopefully connects the complex differential geometry of bubble (tree) to our algebro-geometric or non-archimedean framework.
@@
Under the assumption (3), for $a\in\mathbb{R}_{>0}$, we define $v^{\rm DG}_{a}\colon\mathcal{O}_{\mathcal{X},x}\to\mathbb{R}_{\geq 0}\cup\{+\infty\}$ as follows.  [^16] We do not expect this value would take $\infty$. For $G\in\mathcal{O}_{\mathcal{X},x}$, we set
@@
The following conjecture for the minimum bubble refines the inspiring predictions and discussions in [Sun23, dBS23].
Conjecture 2.30 (cf., [Sun23, 5.4], [dBS23, §4.1]).
Under the Setup 1 with a fixed closed base point $x\in X$, we now again take a quasi-monomial valuation $v_{X}$ of $\mathcal{O}_{X,x}$ which minimizes $\widehat{\operatorname{vol}}_{X}(-)$ ([DS17, §3.2], [Li19, Blu18]) and apply Proposition 2.26 and Definition 2.29 to consider $v_{a}$ and $v_{a}^{\rm DG}$. Then the following holds:
@@
    • $a\in(0,a_{\rm min}]$, $A_{(\mathcal{X},X)}(v_{a})$ stays constant while
@@
    Below, we set $v_{a_{\rm min}}=:v_{\rm min}$ as a notation.  [^17] Note that this is a valuation of the stalk of the total space, unlike $v_{X}$. If $r=1$, then this should be the minimal critical valuation (Theorem 2.4 ($\mathrm{x}$)).
@@
    If $v_{\rm min}$ is divisorial, the birational map $\mathcal{X}^{\prime}_{\rm min}\dashrightarrow\mathcal{X}$ of Theorem 2.4 corresponds to $v_{\rm min}$. Further, it is obtained as an open subset of a plt blow up ([Sho96, 3.1], [Pro00, 2.1]) of $\mathcal{X}$ (e.g., often they are open subsets of weighted blow ups of $\mathcal{X}$ as [dBS23, Remark 11]).
  5. ($\mathrm{v}$)
    If we define ${\rm gr}_{v_{\rm min}}(\mathcal{O}_{\mathcal{X},x}):=\oplus_{m\in{\rm Im}(v_{\rm min})}\mathcal{J}_{m}(v_{\rm min})/\mathcal{J}_{>m}(v_{\rm min})$, then it is a finitely generated $\mathbb{C}[t]$-algebra and $\mathcal{W}(v_{\rm min}):={\rm Spec}_{\mathbb{C}[t]}({\rm gr}_{v_{\rm min}}(\mathcal{O}_{\mathcal{X},x}))$ is a $\mathbb{C}^{*}$-equivariant flat family over $\mathbb{A}^{1}_{t}$ i.e., an affine test configuration. Further, for any $\mathbb{G}_{m}(\mathbb{C})$-invariant section $\sigma\colon\mathbb{A}^{1}_{t}\to\mathcal{W}(v)$, $\sigma(0)$ is always the same which we denote as $p(v)\in\mathcal{W}(v_{\rm min})|_{t=0}$.
  6. ($\mathrm{vi}$)
    There is an isotrivial degeneration181818with a different coordinate/direction from $t$, hence giving test configuration for each fiber of $\mathcal{W}(v_{\rm min})\to\mathbb{A}^{1}_{t}$ $\mathcal{W}(v_{\rm min})\rightsquigarrow\mathcal{B}(v_{\rm min})$ i.e., to be more precise, of $\mathcal{W}(v_{\rm min})\to\mathbb{A}^{1}_{t}$ to another test configuration $\mathcal{B}(v_{\rm min})\to\mathbb{A}^{1}_{t}$ (whose general fiber $\mathcal{B}(v_{\rm min})|_{t=c(\neq 0)}$ is the bubble $B(v)$ in Conjecture 3.1). The limit of $p(v_{\rm min})\in\mathcal{W}(v)|_{t=0}$ is denoted as $q_{\rm min}(v)\in\mathcal{B}(v)|_{t=0}$.
  7. ($\mathrm{vii}$)
    $p(v_{\rm min})\in\mathcal{W}(v_{\rm min})|_{0}$ (resp., $q(v_{\rm min})\in\mathcal{W}(v_{\rm min})|_{0}$) with the natural algebraic torus action, which is associated to the graded structure of ${\rm gr}_{v_{\rm min}}(\mathcal{O}_{\mathcal{X},x})$ and isotrivial degeneration structure of $\mathcal{W}$ and $\mathcal{B}$ give the $2$-step degeneration of [DS14, Li19, LLX20] to the metric tangent cone of $x\in X$.
2.6. A family version of local normalized volume
We analyze the following family version of the normalized local volumes (see the review of the original version in §2.1) motivated by a discussion with C.Spotti (cf., [dBS23, §4.1, esp. after Remark 10]) to whom the author is grateful. On the other hand, we do not mean this subsection to be the main focus of this paper; rather, it is a supplementary for our purpose and arose alongside Conjecture 2.30. ($\mathrm{i}$).
Definition 2.31.
In the above setup,
@@ 2.6. A family version of local normalized volume
    for each $v\in\mathop{\mathrm{Val}}\nolimits(x\in\mathcal{X})$, we define
    $\operatorname{vol}_{\mathcal{X}}(v):=\limsup_{m\to\infty}\dfrac{{\rm length}(\mathcal{O}_{\mathcal{X},x}/\mathcal{J}_{m}(v)|_{X})}{m^{n}/n!},$
    after [ELS03].
@@
    For the same $v$, we also define a normalized version:
    $\widetilde{\operatorname{vol}}_{\mathcal{X}}(v):=(A_{(\mathcal{X},X)}(v))^{n}\cdot\operatorname{vol}_{\mathcal{X}}(v).$
  3. ($\mathrm{iii}$)
    Here is its sibling:
    $\widetilde{\operatorname{vol}}(\mathcal{J}):=(\mathop{\mathrm{lct}}\nolimits((\mathcal{X},X);\mathcal{J}))^{n}\cdot\operatorname{mult}(\mathcal{J}|_{X}),$
    where $\mathcal{J}$ is an arbitrary coherent ideal of $\mathcal{O}_{\mathcal{X}}$ with cosupport $x\in\mathcal{X}$, and $\mathop{\mathrm{lct}}\nolimits$ stands for the log canonical threshold and $\operatorname{mult}$ stands for the multiplicity (here on the level of $X$).
  4. ($\mathrm{iv}$)
    More generally, if we have a graded sequence of coherent ideals $\mathcal{J}_{\bullet}=\{\mathcal{J}_{m}\}_{m=1,2,\cdots}$ with cosupport $x\in\mathcal{X}$ i.e., satisfying $\mathcal{J}_{m_{1}}\mathcal{J}_{m_{2}}\subset\mathcal{J}_{m_{1}+m_{2}}$ for any $m_{1},m_{2}\in\mathbb{Z}_{>0}$ and $\mathcal{J}_{m}\supset\mathfrak{m}_{\mathcal{X},x}^{N_{m}}$ if $N_{m}\gg 0$ for each $m>0$, we similarly define
    $\widetilde{\operatorname{vol}}_{\mathcal{X}}(\mathcal{J}_{\bullet}):=(\mathop{\mathrm{lct}}\nolimits((\mathcal{X},X);\mathcal{J}))^{n}\cdot\operatorname{mult}(\mathcal{J}_{\bullet}|_{X}).$
A subtle but crucial difference here with [Li19, Liu18] is that both terms of $A_{(\mathcal{X},X)}(v)$ and $\mathop{\mathrm{lct}}\nolimits((\mathcal{X},X);\mathcal{J})$ in the above definition reflect the geometry of total space or the process of degeneration, not determined only by the central fiber $X\ni x$. We also note that for any real positive number $c>0$, $\widetilde{\operatorname{vol}}(cv)=\widetilde{\operatorname{vol}}(v)$ by the normalization effect.
Proposition 2.32.
In the above setup, we have
(9) $\inf_{c_{\mathcal{X}}(v)=x}\widetilde{\operatorname{vol}}_{\mathcal{X}}(v)$ $=\inf_{\mathcal{J}\subset\mathcal{O}_{\mathcal{X},x}}\widetilde{\operatorname{vol}}_{\mathcal{X}}(\mathcal{J})$
(10) $=\inf_{\begin{subarray}{c}\mathcal{J}_{\bullet}=\{\mathcal{J}_{m}\}_{m},\\ \mathcal{J}_{m}\subset\mathcal{O}_{\mathcal{X},x}\end{subarray}}\widetilde{\operatorname{vol}}_{\mathcal{X}}(\mathcal{J}_{\bullet})$
(11) $\geq\widehat{\operatorname{vol}}(x\in X).$
We call the quantity (9)$=$(10) as $\widetilde{\operatorname{vol}}_{\mathcal{X}}(x)$. We conjecture (11) to be the equality, where $\widehat{\operatorname{vol}}$ is in the sense of [Li19].
Proof.
It naturally follows from similar arguments to [Liu18, §4.1] but we sketch the proof. The notion of valuation ideal gives a central tool for the proof.
First, for a given coherent ideal $\mathcal{J}$, from the definition of $\mathop{\mathrm{lct}}\nolimits$, there is a divisorial valuation $v_{0}$ with $\mathop{\mathrm{lct}}\nolimits_{(\mathcal{X},X)}(\mathcal{J})=\frac{A_{(\mathcal{X},X)}(v_{0})}{v_{0}(\mathcal{J})}$. For simplicity, we rescale the valuation to make $v_{0}(\mathcal{J})=1$. Then we have $\operatorname{mult}(\mathcal{J}|_{X})\geq\operatorname{vol}(\mathcal{J}_{v}|_{X})$ since $\mathcal{J}^{m}\subset\{f\in\mathcal{O}_{\mathcal{X},x}\mid v_{0}(f)\geq m\}$. Hence, we have $\inf_{c_{\mathcal{X}}(v)=x}\widetilde{\operatorname{vol}}_{\mathcal{X}}(v)\leq\inf_{\mathcal{J}\subset\mathcal{O}_{\mathcal{X},x}}\widetilde{\operatorname{vol}}_{\mathcal{X}}(\mathcal{J})$. For (10), from the arguments in [JM12] and [ELS03], we have $\mathop{\mathrm{lct}}\nolimits((\mathcal{X},X);\mathcal{J}_{\bullet})=\lim_{m\to\infty}m\mathop{\mathrm{lct}}\nolimits((\mathcal{X},X);\mathcal{J}_{m})$ so that
$\inf_{\mathcal{J}\subset\mathcal{O}_{\mathcal{X},x}}\widetilde{\operatorname{vol}}_{\mathcal{X}}(\mathcal{J})\leq\inf_{\begin{subarray}{c}\mathcal{J}_{\bullet}=\{\mathcal{J}_{m}\}_{m},\\ \mathcal{J}_{m}\subset\mathcal{O}_{\mathcal{X},x}\end{subarray}}\widetilde{\operatorname{vol}}_{\mathcal{X}}(\mathcal{J}_{\bullet}).$
Further, for each $v$, the construction of the valuation ideals $\mathcal{J}_{\bullet}=\{\mathcal{J}_{m}(v)\}_{m}$ easily implies
$\inf_{\begin{subarray}{c}\mathcal{J}_{\bullet}=\{\mathcal{J}_{m}\}_{m},\\ \mathcal{J}_{m}\subset\mathcal{O}_{\mathcal{X},x}\end{subarray}}\widetilde{\operatorname{vol}}_{\mathcal{X}}(\mathcal{J}_{\bullet})\leq\inf_{c_{\mathcal{X}}(v)=x}\widetilde{\operatorname{vol}}_{\mathcal{X}}(v).$
Finally, note
$\mathop{\mathrm{lct}}\nolimits(\mathcal{X},\mathcal{J}|_{\mathcal{X}_{0}})\leq\mathop{\mathrm{lct}}\nolimits((\mathcal{X},\mathcal{X}_{0});\mathcal{J}),$
which follow from the inversion of adjunction or comparison of the adjoint ideal to the multiplier ideal (cf., e.g., [Eis11]), that matches to simple consequences of the normal volume comparison in Riemannian geometry under the control of Ricci curvatures from both sides. ∎
We consider the space of real valuations of $\mathcal{O}_{\mathcal{X},x}$ with the center $c_{\mathcal{X}}(v)=x$ and $v(t)=1$ and denote it by $\mathop{\mathrm{Val}}\nolimits^{t\mapsto 1}(x\in\mathcal{X})$. We denote its subset of quasi-monomial valuations as $\mathop{\mathrm{Val}}\nolimits^{\rm qm,t\mapsto 1}(x\in\mathcal{X})$. If $\mathcal{X}$ is associated with a variety $\mathcal{X}_{\eta}$ over convergent power series ring $\mathbb{C}[[t]]^{\rm conv}(\subset\mathbb{C}[[t]])$, $\mathop{\mathrm{Val}}\nolimits^{t\mapsto 1}(x\in\mathcal{X})$ is an open subset of its Berkovich analytification $\mathcal{X}_{\eta}^{\rm an}$ over the fraction field ${\rm Frac}(\mathbb{C}[[t]]^{\rm conv})=\mathbb{C}((t))^{\rm conv}$, so we have the weak topology on it. A subtle caution here is such valuation in $\mathop{\mathrm{Val}}\nolimits^{t\mapsto 1}(x\in\mathcal{X})$ does not necessarily restrict to (unique) valuation at $x\in X$ in general.
Example 2.33 (Equisingular case).
If $\mathcal{X}=X\times\Delta$, it is natural to consider $\mathbb{C}^{*}$-invariant valuation i.e., valuation $v$ of the form $v(\sum_{i}f_{i}t^{i})=\min_{i}\{v_{X}(f_{i})+iw\}$ for some valuation $v_{X}$ of $\mathcal{O}_{X,x}$ with the center $\mathfrak{m}_{X,x}$ and a positive number $w$. Then it follows from a standard calculation to see $A_{(\mathcal{X},x)}(v)=A_{X}(v_{X})$, $\operatorname{vol}_{\mathcal{X}}(\mathcal{J}_{m}(v))=\operatorname{vol}(v_{X})$ so that $\widetilde{\operatorname{vol}}_{\mathcal{X}}(v)=\widehat{\operatorname{vol}}_{X}(v_{X})$. Hence the problem of minimizing $\widetilde{\operatorname{vol}}$ should be reduced to $\widehat{\operatorname{vol}}$ in the sense of [Li19].
3. Discussions on more general bubbles
3.1. Deeper bubbles for klt degenerations
For the structure formed by more general bubbles, we have rougher predictions. For convenience of discussions, we implicitly take a rescaling normalization of valuations as $v(t)=1$ in this section i.e., multiply $a$ to those considered in section 2.5.
Conjecture 3.1.
We fix an affine finite type log-terminal faithfully flat morphism $\pi\colon\mathcal{X}\to\Delta^{\rm conv}(\mathbb{C})$, where $\Delta^{\rm conv}(\mathbb{C})$ denotes the convergent spectrum $\mathop{\mathrm{Spec}}\nolimits(\mathbb{C}[[t]]^{\rm conv}:=\mathcal{O}_{\mathbb{A}^{1},0}^{\rm an})$, with $v(t)=1$. (Note that in particular, we have its restrictions of the form $\mathcal{Y}\to\Delta_{\epsilon}:=\{t\in\mathbb{C}\mid|t|<\epsilon\}$ for some $\epsilon>0$ with the same central fibers.) Then the following holds.
  1. ($\mathrm{i}$)
    Consider the rescaling equivalence classes of non-flat bubbles at $x\in\mathcal{X}$, which we write $\mathfrak{B}(x\in\mathcal{X})=\{(p\in B,g)\}/\mathbb{R}_{>0}$, and the class $(p\in B,g)$ represents as $[p\in B,g]$. There is a bijective correspondence between $\mathfrak{B}(x\in\mathcal{X})=\{(p\in B,g)\}/\mathbb{R}_{>0}$ and a subset $\mathop{\mathrm{Crit}}\nolimits(x\in\mathcal{X})\subset\mathop{\mathrm{Val}}\nolimits^{t\mapsto 1}(x\in\mathcal{X})$. We write the latter as $\mathop{\mathrm{Crit}}\nolimits(x\in\mathcal{X})$ and the bijective maps as
    $\nu\colon$ $\mathfrak{B}(x\in\mathcal{X})\xrightarrow{\,\smash{\raisebox{-1.50696pt}{$\scriptstyle\simeq$}}\,}\mathop{\mathrm{Crit}}\nolimits(x\in X)\subset\mathop{\mathrm{Val}}\nolimits^{t\mapsto 1}(x\in\mathcal{X}),$
    which are inverse to each other. In particular, the set of bubbles is embedded into a non-archimedean Berkovich analytic space.
  2. ($\mathrm{ii}$)
@@ 3.1. Deeper bubbles for klt degenerations
    $\mathop{\mathrm{Crit}}\nolimits(x\in\mathcal{X})\subset\mathop{\mathrm{Val}}\nolimits^{t\mapsto 1}(x\in\mathcal{X})$ and compatibly the moduli of bubble classes $\mathfrak{B}(x\in\mathcal{X})$ has a locally PL structure.
@@
To make sense of the item ($\mathrm{ii}$), we prepare the following after the general ideas in [Berk99, KS06, BFJ08, BFJ16].

### v2

@@ 2.5. Valuative approach
Taking Theorem 2.4 ($\mathrm{x}$) and Proposition 2.24 ($\mathrm{ii}$) into account, for instance, it is reasonable to pass our context to general valuations and discuss characterization of the critical valuations in the sense of Theorem 2.4 ($\mathrm{x}$). Our statements in this subsection are totally coordinates-free.
@@
    For each $v_{X}\in\mathop{\mathrm{Val}}\nolimits^{\rm qm}(x\in X)$, $a\in\mathbb{R}_{>0}$, with $a\ll 1$, say $a\in(0,a_{0})$ for some $a_{0}\in\mathbb{R}_{>0}$, we have a unique valuation $v\in\mathop{\mathrm{Val}}\nolimits^{\rm qm}(x\in\mathcal{X})$ with $\mathcal{J}_{m}(v_{a})|_{X}=J_{m}(v_{X})$ for all $m$ (or equivalently, for all $m\gg 0$) and $v_{a}(t)=\frac{1}{a}$.
@@
The reason of our notation of the weight as reciprocal $\frac{1}{a}$ comes from our interest in $a=0$ case as initial point (cf., also Remark 2.28).
@@
If $v_{a}(t)=\frac{1}{a}>v_{a}(z_{i})$, then by (14), the independencity of $v_{a}$ with respect to the lifted coordinates $\tilde{z_{i}}$ easily follows. This confirms ($\mathrm{i}$). On the other hand, if we fix a section $\sigma$, then we can impose the condition $\tilde{z_{i}}(\sigma(t))=0$ for each $i$. Under such constraints, $f_{i}$ of (14) does not have constant terms i.e., positive multiplicities with respect to $\tilde{z_{i}}$s. In that case, if $v_{a}(t)=\frac{1}{a}>\max\{c_{i}\}-\min\{c_{i}\}$, then the obtained quasi-monomial $v_{a}$ does not depend on the lift $\tilde{z_{i}}$s. Hence the proof of ($\mathrm{ii}$). ∎
@@
S.Boucksom kindly suggested to compare the construction with the composite (cf., e.g., [FK18, §6.4]) of the discrete valuation ${\rm ord}_{\mathcal{X}_{0}}$ on $\mathcal{O}_{\mathcal{X},x}$ and $v_{X}$ on $\mathcal{O}_{\mathcal{X}_{0},x}$. Although it gives a height $2$ (not real-valued) valuation, the author sees it as a kind of infinitesimal analogue of the above construction $\{v_{a}\}_{a}$, which may be its integral curve in some unknown sense.
@@
Suppose $\mathcal{X}$ has a simultaneous log resolution, such as $2$-dimensional ADE singularity case. Then, we can apply the above proof to that simultaneous resolution so that the central fiber of $\tilde{X}$ is smooth (irreducible) and $\tilde{z}_{i}$ are all different from $t$, so that we see $a_{0}=\infty$. On the other hand [dBS24, Example 7] infers that $a_{0}$ could be finite in general, for $n(>2)$-dimensional $A$-type singularities.
In this paper, we introduce yet another couple of “Yau-Tian-Donaldson type” conjectures for mild degenerations, after [Sun23, dBS24]. It hopefully connects the complex differential geometry of bubble (tree) to our algebro-geometric or non-archimedean framework.
@@
Under the assumption (3), for $a\in\mathbb{R}_{>0}$, we define $v^{\rm DG}_{a}\colon\mathcal{O}_{\mathcal{X},x}\setminus\{0\}\to\mathbb{R}_{\geq 0}\cup\{+\infty\}$  [^17] We do not expect this value can take $\infty$ in reality. as follows. For $G\in\mathcal{O}_{\mathcal{X},x}$, we set
@@
The following conjecture for the minimum bubble refines the inspiring predictions and discussions in [Sun23, dBS24] via valuations, especially by variation of log discrepancies and the above type valuation (Definition (2.30)).
Conjecture 2.31.
Under the Setup 1 with a fixed closed base point $x\in X$, we now again take the quasi-monomial valuation $v_{X}$ of $\mathcal{O}_{X,x}$ which minimizes $\widehat{\operatorname{vol}}_{X}(-)$ ([DS17, §3.2], [Li19, Blu18]) and apply Proposition 2.27 and Definition 2.30 to consider $v_{a}$ and $v_{a}^{\rm DG}$. Then the following holds:
@@
    • The general fibers of a test configuration ${\rm Spec}({\rm gr}_{v_{\rm min}}\mathcal{O}_{\mathcal{X},x})\to\mathbb{A}^{1}$ is the semistable minimal bubble $\mathcal{X}^{\prime}_{{\rm min},0}$. (compare Corollary (2.26)). In relation with it, if $r=1$, then $v_{\rm min}$ should be the minimal critical valuation (Theorem 2.4 ($\mathrm{x}$)).
    • $a\in(0,a_{\rm min}]$, $A_{(\mathcal{X},X)}(v_{a})$ stays constant (cf., Proposition 2.32 below) while
@@
    (family version of local normalized volume) For an arbitray valuation $v$ of $\mathcal{O}_{x,\mathcal{X}}$ with center $x\in\mathcal{X}$, we consider $\widetilde{\operatorname{vol}}_{\mathcal{X}}(v):=(A_{(\mathcal{X},X)}(v))^{n}\cdot\operatorname{vol}_{\mathcal{X}}(v).$ Then, we have  [^19] In the version 1 on arXiv, we include the proof of $\geq$ direction and other related statements. We temporarily cut the details just to make the exposition more concise with clearer focus. $\inf_{c_{\mathcal{X}}(v)=x}\widetilde{\operatorname{vol}}_{\mathcal{X}}(v)=\widehat{\operatorname{vol}}(x\in X).$
The above conjectures are obtained after experimental calculations of various examples, including non-degenerate hypersurfaces. We partially confirm it as follows.
Theorem 2.32.
Under the same setup, Conjecture 2.31 is partially true in the following sense: there is a positive real number $a_{\rm min}$ and a family of quasi-monomial valuations $v_{a}$ of $\mathcal{O}_{\mathcal{X},x}$ ($0<a\leq a_{\rm min}$) such that the first and second bulleted items in ($\mathrm{i}$) hold.
  • (compare Proposition 2.27) The valuative ideals satisfy $\mathcal{J}_{m}(v_{a})|_{X}=J_{m}(v_{X})$ for all $m$ and $v_{a}(t)=\frac{1}{a}$.
  • The general fibers of a test configuration ${\rm Spec}({\rm gr}_{v_{a_{\rm min}}}\mathcal{O}_{\mathcal{X},x})\to\mathbb{A}^{1}$ is the semistable minimal bubble $\mathcal{X}^{\prime}_{{\rm min},0}$. (compare Corollary (2.26)).
  • $A_{(\mathcal{X},X)}(v_{a})$ stays constant for $0<a\leq a_{\rm min}$ (cf., Proposition 2.32 below)
Note that the last item could be roughly regarded as a difficult version of adjunction, once we have $\mathcal{J}_{m}(v_{a})|_{X}=J_{m}(v_{X})$ for all $m$, but is much harder and require $a<a_{\rm min}.$
Proof.
The following depends on the arguments used in previous (sub)sections and [Od24b, §3]. As before, we take the volume-minimizing valuation $v_{X}$, realize it by weights on the coordinates of local embedding of $X$ to affine spaces, consider the corresponding Reeb vector $\xi\in N\otimes\mathbb{R}$ to $v_{X}$ and replace it by an (integral) element of $N$ whose direction approximates the original $\mathbb{R}_{>0}\xi$ i.e., in $\tau\cap N$ in the notation of [Od24b, §3]. It is easy to see that assertions of the above proposition for such replacement $\xi$ implies the same for the original $\xi$.
Recall that the proof of Theorem 2.4 shows that the condition of Lemma 2.25 is satisfied for certain $w_{0}\in\mathbb{Z}_{>0}$, after replacing $\xi$ by its positive integer multiple, so that one can apply it. If we set $a_{\rm min}:=\frac{1}{w_{0}}$, Lemma 2.25 (i) shows that the filtration on $\mathcal{O}_{\mathcal{X},x}$ with respect to the weights $(a\xi,1)$ with $a<a_{\rm min}$ gives a valuation on $\mathcal{O}_{\mathcal{X},x}$, which we denote as $av_{a}$. This is because the spectrum of the graded ring is still isomorphic to the K-semistable Fano cone $W$ after the perturbation of $\xi$ (see e.g., [LX18, 2.10]) and in particular integral. This gives the valuation corresponding to the exceptional divisor of the weighted blow up with the weights $(a\xi,1)$. Further, it follows from the construction that $\mathcal{J}_{m}(v_{a})|_{X}=J_{m}(v_{X})$ for all $m$ and $v_{a}(t)=\frac{1}{a}$.
Below, we show that $A_{(\mathcal{X},X)}(v_{a})$ does not depend on $a$ which are integers. By rescaling $v_{X}$ and base change of $\mathcal{X}\to\Delta$ it implies the constancy of $A_{(\mathcal{X},X)}(v_{a})$ for all rational $a<a_{\rm min}$ hence the same for all real $a<a_{\rm min}$ by its continuity with respect to $a$. Therefore, we can and do assume $a(<a_{\rm min})$ is positive integer and only need to prove the independency of $A_{(\mathcal{X},X)}(v_{a})$. Following Lemma 2.25 (i) we consider the corresponding normalized weighted blow up of $\mathcal{X}$ for the same weights $(a\xi,1)$ which we denote as $\tilde{\rm Bl}_{(a\xi,1)}(\mathcal{X})(\to\mathcal{X})\to\Delta$. Its central fiber consists of two irreducible components i.e., the strict transform $X^{\prime}$ of $X$ and a compactification of $W$. Note that since the latter is the exceptional divisor and the total central fiber is principal divisor of $\tilde{\rm Bl}_{(a\xi,1)}(\mathcal{X})$, both components are $\mathbb{Q}$-Cartier. Further, from [LX20, 1.3], [LX18, 2.10, 4.7], one can and do assume that our approximant $\xi$ corresponds to the exceptional divisor $e$ of a plt blow up of $X$ i.e., ${\rm val}_{\xi}=c{\rm ord}_{e}$ as valuations on $\mathcal{O}_{X,x}$ with some positive real constant $c$. Then, for each $\mathbb{Z}_{>0}\ni a(<a_{\rm min})$, we have
$A_{(\mathcal{X},X)}(v_{a})$ $=\frac{1}{a}A_{(\mathcal{X},X)}({\rm ord}_{W})$
$=\frac{1}{a}A_{X}({\rm ord}_{W}|_{X})$
$=A_{X}(v_{X})=A_{X}(c{\rm ord}_{e}),$
which does not depend on $a$. Thus we confirm the third bulleted item. Also, Lemma 2.25 (ii) implies that the general fibers of ${\rm Spec}({\rm gr}_{v_{a_{\rm min}}}\mathcal{O}_{\mathcal{X},x})\to\mathbb{A}^{1}$ is the semistable minimal bubble $\mathcal{X}^{\prime}_{{\rm min},0}$. We complete the proof of the second bulleted item as well. ∎
3. Discussions on more general bubbles
This last short section takes a few pages to discuss conjectural pictures for future direction. In the first subsection §3.1, we still discuss klt degenerations as in Setup 1 (“type I”, but essentially polarized) and speculate on the whole structure formed by deeper bubbles. Then §3.2 reviews our speculations on the structure of bubblings for so-called maximal degeneration case or intermediate case (“type II” or higher), following [Od22a, Od22b].
3.1. Deeper bubbles for klt degenerations
For the structure formed by all the bubbles for the same klt type degenerations as Setup 1, we only have rougher brave predictions as follows. For convenience of discussions, we implicitly take a rescaling normalization of valuations as $v(t)=1$ in this section i.e., multiply $a$ to those considered in section 2.5.
Conjecture 3.1.
We fix an affine finite type log-terminal faithfully flat morphism $\pi\colon\mathcal{X}\to\Delta^{\rm conv}(\mathbb{C})$, where $\Delta^{\rm conv}(\mathbb{C})$ denotes the convergent spectrum $\mathop{\mathrm{Spec}}\nolimits(\mathbb{C}[[t]]^{\rm conv}:=\mathcal{O}_{\mathbb{A}^{1},0}^{\rm an})$, with $v(t)=1$. (Note that in particular, we have its restrictions of the form $\mathcal{Y}\to\Delta_{\epsilon}:=\{t\in\mathbb{C}\mid|t|<\epsilon\}$ for some $\epsilon>0$ with the same central fibers.) We denote the generic fiber as $\mathcal{X}_{\eta}$ and its Berkovich analytification as $\mathcal{X}_{\eta}^{\rm an}$ Then the following holds.
@@ 3.1. Deeper bubbles for klt degenerations
    Consider the rescaling equivalence classes of non-flat bubbles at $x\in\mathcal{X}$, which we write $\mathfrak{B}(x\in\mathcal{X})=\{(p\in B,g)\}/\mathbb{R}_{>0}$, and the class $(p\in B,g)$ represents as $[p\in B,g]$. There is a bijective correspondence between $\mathfrak{B}(x\in\mathcal{X})=\{(p\in B,g)\}/\mathbb{R}_{>0}$ and a set of valuations $\mathop{\mathrm{Crit}}\nolimits(x\in\mathcal{X})\subset\mathcal{X}_{\eta}^{\rm an}$. We write the latter as $\mathop{\mathrm{Crit}}\nolimits(x\in\mathcal{X})$ and the bijective maps as
    $\nu\colon$ $\mathfrak{B}(x\in\mathcal{X})\xrightarrow{\,\smash{\raisebox{-1.50696pt}{$\scriptstyle\simeq$}}\,}\mathop{\mathrm{Crit}}\nolimits(x\in X)\subset\mathcal{X}_{\eta}^{\rm an},$
    which are inverse to each other. In particular, the set of bubbles (with both the algebraic structures and metrics) has an natural embedding into a non-archimedean Berkovich analytic space.
@@
    $\mathop{\mathrm{Crit}}\nolimits(x\in\mathcal{X})\subset\mathcal{X}_{\eta}^{\rm an}$ and compatibly the moduli of bubble classes $\mathfrak{B}(x\in\mathcal{X})$ has a locally PL structure.
@@
To make sense of the item ($\mathrm{ii}$) in general, we prepare the following after the general ideas in [Berk99, KS06, BFJ08, BFJ16] for the case $\mathcal{X}_{\eta}$ is smooth.